9. Find the 10th degree Taylor polynomial for the following functions: a) cos(x2) b) sin(2x) c) ex+1 d) ex? cos(x3)

c,d,e please, clean work, clean handwriting please
Answer #1

© ex+1, The taylor polynomial is given as f +(-) = P(W) = 2 070)_ OX f (x) ~ Pla K=0 KI k- ok! so we need to find the derivatè cosx) = f(i) The results of desvatives for this Bunction are to wide, the reture. I only write Here all the desivatives. PlSinca?).fo., {CO) - - 2x(85963?) + (-?-1) 2013 (32) 2 (x+12 2 (a) = – 2 [(2n + ant - ²+)-sinlan?) + (3 nt+1-1) Cosa3] It az